OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (slaaibak):

If lines are perpendicular, m1 * m2 = -1 That's how you will find the slope of the line. Once you have the slope, use this formula: \[y - y_1 = m(x - x_1)\]

OpenStudy (anonymous):

But isn't the slope intercept formula y = mx + b?

OpenStudy (slaaibak):

Once you plug in the values m, y1 and x1, you will be able to write it in the form of y = mx + b

OpenStudy (anonymous):

Ah, well thank you then!

OpenStudy (slaaibak):

I'll demonstrate: Slope of the line you need to find = 1/4, since -4 * 1/4 = -1 Now, you have m=1/4 you have the point (-16, -11) Using my formula: y - (-11) = 1/4 (x - (-16)) y + 11 = 1/4x + 4 y=1/4x - 7

OpenStudy (anonymous):

Okay, thank you. That makes a lot more sense c:

OpenStudy (anonymous):

But where did the 1/4 come from exactly?

OpenStudy (slaaibak):

The lines are perpendicular. slope of line1 * slope of line 2 = -1 That's a rule for lines to be perpendicular. now you have the slope of line 1: -4 * m2 = -1 m2 = -1/-4 = 1/4

OpenStudy (anonymous):

Oh okay, that makes sense. Thank you.